Tha area of a rectangle is 108 cm squared. The ratio of the width to the lenth is 3:4. Find the length and the width.

Answer by checkley77(12844)

3:4 or 3n+4n
area=108
3n*4n=108
12n^2=108
n^2=108/12
n^2=9
n=sqrt9
n=3
3*3=9 ans. for one side.
4*3=12 ans. for the other side.
Proof:
9812=108